how many liters of 20% salt solution must be added to 72 liters of 59% salt solution to get a 44% salt solution?
Answer by alicealc(293)

volume of 20% salt solution = x

20% * x + 59% * 72 = 44% * (x + 72)
0.2x + 0.59 * 72 = 0.44 * (x + 72)
0.2x + 42.48 = 0.44x + 0.44 * 72
0.2x + 42.48 = 0.44x + 31.68
42.48 - 31.68 =0.44x - 0.2x
10.8 = 0.24x
x = 10.8/0.24 = 45

so, the volume of 20% salt solution to be added is 45 liters